# Capacitance-based Fermion parity read-out and predicted Rabi oscillations in a Majorana nanowire

### v2

@@ Abstract
The left Majorana nanowire in Fig. 1 contains a QD tunnel-coupled to both ends, enclosing a flux $\Phi$. The QD enables tunneling via elastic co-tunneling [15], including cross-Andreev processes [16, 17, 10], which effectively act as a direct tunneling amplitude between the nanowire ends [8]. Multi-particle excitations [15] generated in the co-tunneling process are virtual and only renormalize the tunneling amplitude. This tunneling is modulated by the gate voltage $V_{QD}$, also used for capacitance measurements. We model the QD as a single non-interacting orbital with energy level $V_{QD}$, adjusting tunneling to match capacitance data. This allows parity measurement using a non-interacting BdG Hamiltonian for the Majorana nanowire [18] (see Sec. $S_{1}$ in SI for details). The magnetic flux $\phi$ is in units where the Hamiltonian periodicity matches the flux quantum $hc/e=2$. Topological stability ($T_{S}$) is defined as the class D topological invariant [1] multiplied by the ratio of the lowest eigenvalues of the Hamiltonian (without the QD) at $\phi=0$ and $\phi=1$, helping assess wire topology and length effects. The QD hopping efficiency is set to $\lambda_{QD}=0.03$ to match energy splittings in the tens of mK, consistent with experiments. For disordered wires, the disorder potential follows a Gaussian distribution with correlation length $\xi_{dis}\sim 30nm$ and root mean square amplitude $V_{RMS}$.
@@
where $\Gamma_{C}=2\alpha^{2}e^{2}/k_{B}=1.9\,fF-K$ is a constant that converts the capacitance $C(\omega)$ to femtofarads (fF) when the energies $E_{n}$ are in units of Kelvin (K) after adding a lever arm $\alpha=2^{-1/2}$ that would be relevant to recent experiment [8]. In the above |$\ket{E_{n}}$ are eigenstates of the BdG Hamiltonian of the system, and $\tau_{z}$ is the matrix coupling to charge density, both written in the Nambu basis (see SI Sec.$S_{1}$ for details). The read-out scheme used in the recent experiments [8] involves studying the dependence of $C$ on the magnetic flux $\phi$ as well as the parity of the wire, which can be changed by switching the occupancy of the state of the smallest energy states from 00 to $1$ and vice-versa.
Figure 2: (a) Wave-function of the lowest energy eigenstate, (b) energy in K and (c) capacitance for the occupied and unoccupied state as a function of magnetic flux $\phi$ (in units where $hc/e=2$) for an $L=3\,\mu m$ Majorana nanowire without disorder at $\mu=0\,K$, $B_{x}=0.28\,T$ and $V_{QD}=0.5\,K$. The wave-functions for this ideal case show clear presence of end MZMs with a topological stability [1] (see Sec. $S_{1}$ of SI for details) $T_{S}=-0.99$ and MZM splitting for the isolated wire $E_{M}=30\,\mu K$. As expected [8] we see a shift of $\Delta\phi=1$ between the even and odd parity sectors. The panels (d,e,f) are the corresponding result for the Majorana nanowire in a non-topological phase with $T_{S}=0.49$. The parameters for the phase are $B_{x}=0.4\,T$, $\mu=5.5\,K$, $V_{QD}=5.5\,K$ and $L=3\,\mu m$. The low energy states are ABSs generated by a potential (see text for details) with parameters $V_{abs}=4.5\,K$ and $L_{abs}=200\,nm$.
@@
While capacitance-based parity readout works well for ideal Majorana nanowires, its effectiveness in regimes [19] with trivial zero-bias peaks from disorder or end potentials remains uncertain. In Fig. 2(d–f), we examine cases where zero modes originate from non-topological near-zero-energy ABSs, often induced by end quantum dot potentials [20, 21, 22, 23, 24]. Fig. 2(d) shows such ABSs forming due to an end quantum well potential (see SI Sec. $S_{1}$). Despite their low energy, Fig. 2(e) indicates that ABS spectra exhibit flux modulation smaller than the average energy, confirming the wire’s non-topological nature. Capacitance readout in Fig. 2(e) reveals a clear parity difference, but the flux dependence is dominated by an $hc/2e$ periodic modulation that flips sign between parities, unlike the ideal Majorana case where the $hc/e$ component maintains its sign. This aligns with recent capacitance analyses based on a minimal low-energy ABS model [8].
@@
Figure 3: (a) Wave-function of the lowest energy eigenstate, (b) energy in K and (c) capacitance for the occupied and unoccupied state as a function of magnetic flux $\phi$ (in units where $hc/e=2$) for an $L=3\,\mu m$ Majorana nanowire in the topological patch regime [11] at $\mu=0.2\,K$ and $B_{x}=0.42\,T$, $V_{QD}=0.2\,K$, disorder strength $V_{RMS}=7.5\,K$. The wave-functions for this ideal case show clear presence of end MZMs with a topological stability [1] (see Sec. $S_{1}$ of SI for details) $T_{S}=-0.621$. As expected [8] we see an approximate shift of $\Delta\phi=1$ between the even and odd parity sectors. The panels (d,e,f) are the corresponding result for the Majorana nanowire in a non-topological phase with $T_{S}=0.13$. The parameters for the phase are $B_{x}=0.56\,T$, $\mu=2.0\,K$, $V_{QD}=2.0\,K$ $V_{RMS}=15\,K$ and $L=3\,\mu m$.
Three-terminal measurements of tunnel barrier dependence can rule out local zero-energy ABSs [25]. However, low-energy states from intermediate disorder, known as topological patches [11, 12], may explain experimentally observed topological regions [25]. Fig. 3(a) shows the lowest-energy eigenstate in a weakly disordered Majorana nanowire, where despite a negative topological stability ($T_{S}$), the ABS wavefunction lacks significant end localization, distinguishing it from true MZMs. Still, Fig. 3(b) reveals an $hc/2e$ flux shift between parity sectors, resembling the MZM spectrum in Fig. 2(b). Similarly, capacitance in this regime (Fig. 3(c)) exhibits a notable $hc/e$ component, flipping sign like in the topological case, though with asymmetric oscillation amplitudes, consistent with some experimental samples [8]. Topological stability in the patch regime [11] is highly disorder-dependent. Stronger disorder (Fig. 3(d)) leads to shorter localization compared to Fig. 3(a). Despite a trivial $T_{S}$ ($T_{S}>0$), Fig. 3(e–f) still shows substantial $hc/e$ components, resembling the topological case in the upper panels.
@@
where $H_{L,R}$ are given by Eq. 3 except that the sites are restricted to $r\neq 0$ together with $\lambda_{QD}=0$ to remove the QD and $\lambda(\tau)$ is the dimensionless tunneling strength between the wires. Note that this set-up differs fundamentally from previous proposals for studying coherence in nanowire qubits [26], which requires inter-wire parity measurements. which requires inter-wire parity measurements. Instead, the set-up in Fig. 1, similar to previous experiments [13], is based on intra-wire parity measurements together with inter-wire tunneling that is set to a non-zero value $\lambda(\tau>0)=\lambda$ after initializing with $\lambda(\tau=0)=0$. Restricting the dynamics to the lowest two positive energy state (one from each nanowire) and considering a specific total Fermion parity sector allows one to map the dynamics using the Bloch-Redfield master equation of a standard two-level spin system [27] that is characterized by a Rabi oscillation frequency $\Omega$, relaxation time $T_{1}$, dephasing time $T_{2}$ as well as an amplitude for the Rabi oscillation of the Fermion parity of the left wire (see SI Sec. $S_{2}$ for details). The decoherence is generated by bath couplings that couple to the gate voltage in each wire with magnitude $\gamma_{L,R}\sim 80\,mK$ and are chosen to reproduce conventional charge qubit dephasing [28] (see SI Sec. $S_{3}$ for details). Inter-wire parity or capacitance measurements, which are necessary to perform topological operations [6], allow one to estimate coherence of the qubits [26] without inter-wire tunneling. The estimate of dephasing rates in this case takes the form of a decoherence time $T_{1}^{\prime}$ for the inter-wire fermion parity. Different choices of ends for the inter-wire fermion parity measurement $X$ and $Y$ are expected to generate different superpositions that differ by a phase shift $\delta\phi$ that can also characterize a topological qubit (see SI Sec. $S_{4}$ for details).

### v3

@@ I Introduction
II Parity read-out
The left Majorana nanowire supporting $ABS_{L}$ shown in Fig. 1 contains a QD tunnel coupled to both ends of the Majorana nanowire with a flux $\Phi$ enclosed by the loop. The essential function of the QD in the experiment [8] is to provide tunneling between the ends of the nanowire through elastic co-tunneling [15], which includes cross-Andreev processes that maybe comparable to the direct tunneling [16, 17, 10]. The Coulomb interaction in the QD is essential to enhance the single particle (charge) gap of approximately $45\,\mu eV$ [8] in the QD beyond the single-particle level spacing of $20\,\mu eV$, which would be quite low for a $2.4\,\mu$m long QD [8]. For weak tunneling that leaves the QD in the strongly Coulomb blockaded regimes [8], we can represent the states in terms of a single electron creation operator $d^{\dagger}$ corresponding to the single-particle level closest to the Fermi level together with a host of lower energy particle-hole excitations in the QD levels [15]. For sufficiently low voltages, frequencies and tunnel coupling, these particle-hole excitations are gapped by the single-particle level spacing so that their contribution to the co-tunneling process is a perturbative renormalization of the tunneling amplitude. In this limit, the coherent elastic co-tunneling across the QD is phenomenologically identical to a tunneling amplitude between the ends of the nanowire. Furthermore, the elastic co-tunneling is modulated by the gate voltage $V_{QD}$ at the QD that is also used to measure the capacitance. We model the low-energy properties of the QD at a phenomenological level by tunneling through a single non-interacting orbital with an energy level at $V_{QD}$. The magnitude of the tunneling is adjusted to match the measured value of capacitance [8]. Potential complications of the multi-level interacting QD physics that not captured by this renormalized parameter model will be discussed in the conclusion.
The model for the QD described above allows us to understand the parity measurement using the configuration of the left Majorana nanowire in Fig. 1 in terms of a non-interacting Bogoliubov de Gennes (BdG) Hamiltonian, $H(\phi)$, that is the Majorana nanowire model [18] (see Appendix. A for details). The magnetic flux $\phi$ in the measuring loop containing the QD in Fig. 1 is in units where the periodicity of the Hamiltonian in $\phi$ i.e. the flux quantum is set to $hc/e=2$. We quantify the degree to which the wire is topological by the topological stability $T_{S}$ defined as the product of the class D topological invariant [1] with the ratio of lowest eigenvalues for the Hamiltonian $H$ (without the QD) with flux $\phi=0$ and $\phi=1$(see Appendix A for details). This metric, apart from measuring the topological invariant (indicated by a negative sign), can indicate a short wire when this ratio differs substantially from unity. Thus a value of $T_{S}$ substantially larger than -1 indicates a poorly defined topological phasse because of finite size effects [11]. The hopping efficiency through the QD (for our numerical results) is chosen to be $\lambda_{QD}=0.03$ so as to obtain energy splittings in the range of several tens of mK (consistent with the current experimental situation). When studying disordered wires, we choose the disorder potential for the Majorana nanowire to be Gaussian distributed with a correlation length of $\xi_{dis}\sim 30nm$ with a root mean square amplitude $V_{RMS}$.
@@ II Parity read-out
where $\Gamma_{C}=2\alpha^{2}e^{2}/k_{B}=1.9\,fF-K$ is a constant that converts the capacitance $C(\omega)$ to femtofarads (fF) when the energies $E_{n}$ are in units of Kelvin (K) after adding a lever arm $\alpha=2^{-1/2}$ that would be relevant to recent experiment [8]. In the above |$\ket{E_{n}}$ are eigenstates of the BdG Hamiltonian of the system, and $\tau_{z}$ is the matrix coupling to charge density, both written in the Nambu basis (see Appendix A for details). The read-out scheme used in the recent experiments [8] involves studying the dependence of $C$ on the magnetic flux $\phi$ as well as the parity of the wire, which can be changed by switching the occupancy of the state of the smallest energy states from 00 to $1$ and vice-versa.
Figure 2: (a) Wave-function of the lowest energy eigenstate, (b) energy in K and (c) capacitance for the occupied and unoccupied state as a function of magnetic flux $\phi$ (in units where $hc/e=2$) for an $L=3\,\mu m$ Majorana nanowire without disorder at $\mu=0\,K$, $B_{x}=0.28\,T$ and $V_{QD}=0.5\,K$. The wave-functions for this ideal case show clear presence of end MZMs with a topological stability [1] (see Appendix A for details) $T_{S}=-0.99$ and MZM splitting for the isolated wire $E_{M}=30\,\mu K$. As expected [8] we see a shift of $\Delta\phi=1$ between the even and odd parity sectors. The panels (d,e,f) are the corresponding result for the Majorana nanowire in a non-topological phase with $T_{S}=0.49$. The parameters for the phase are $B_{x}=0.4\,T$, $\mu=5.5\,K$, $V_{QD}=5.5\,K$ and $L=3\,\mu m$. The low energy states are ABSs generated by a potential (see text for details) with parameters $V_{abs}=4.5\,K$ and $L_{abs}=200\,nm$.
II.1 Parity results
@@
While capacitance-based parity readout works well for ideal Majorana nanowires, its effectiveness in regimes [19] with trivial zero-bias peaks from disorder or end potentials remains uncertain. In Fig. 2(d–f), we examine cases where zero modes originate from non-topological near-zero-energy ABSs, often induced by end quantum dot potentials [20, 21, 22, 23, 24]. Fig. 2(d) shows such ABSs forming due to an end quantum well potential (see Appendix A). Despite their low energy, Fig. 2(e) indicates that ABS spectra exhibit flux modulation smaller than the average energy, confirming the wire’s non-topological nature. Capacitance readout in Fig. 2(e) reveals a clear parity difference, but the flux dependence is dominated by an $hc/2e$ periodic modulation that flips sign between parities, unlike the ideal Majorana case where the $hc/e$ component maintains its sign. This aligns with recent capacitance analyses based on a minimal low-energy ABS model [8].
Note that while the detailed results in Fig. 2 depend on a specific choice of parameters listed in the caption that are used in the Hamiltonian described in Appendix A, the qualitative features are generic for the corresponding state i.e. topological MZM or ABSs. The parameters listed in the caption are in the range that experiments [8] tune within and in this case are fine-tuned to produce a small MZM splitting or ABS end states. The parameter choice is motivated from previous descriptions of the phase diagram of the MZM nanowire model [2, 7] in the field [11, 23] among others and therefore will not be repeated here. The local zero-energy ABS model described above can practically be ruled out by three terminal measurements of the tunnel barrier dependence of the local and non-local conductance [25]. However, fine-tuned low energy states resulting from intermediate strength disorder called topological patches [11, 12] can be a viable explanation for the topological regions seen experimentally [25]. Fig. 3(a) shows the wave-function of the lowest energy eigenstate of the weakly disordered Majorana nanowire with parameters chosen to generate such a zero-energy state following previous work [11] on the model. Despite the topological stability with $T_{S}$ being negative, which would nominally be defined as topological based on the usual topological invariant [1], the lowest energy ABS wave-function has rather small weight at the end and cannot reasonably be construed as being composed of end-localized MZMs. Despite this, Fig. 3(b) shows that the lowest energy state has a flux dependence with a substantial $hc/2e$ shift between different parity sectors, which is rather similar to the topological MZM spectrum seen in Fig. 2(b). Similar to the topological case, the capacitance in the topological patch regime seen in Fig. 3(c) shows a significant $hc/e$ periodic component. Furthermore, this component flips sign similar to the topological case (i.e. Ref. 2(c)). However, unlike the ideal case, the amplitude of oscillations in the different parity sector are different as is seen in some of periods of the capacitance data measured in recent experiments [8]. It should be noted that identifying the oscillation amplitude in the experiment is difficult because of the Aharonov Bohm oscillation amplitude varying between different periods of the oscillation due to magnetic field induced shifts. Topological stability in the patch regime [11] is highly disorder-dependent. Stronger disorder (Fig. 3(d)) leads to shorter localization compared to Fig. 3(a). Despite a trivial $T_{S}$ ($T_{S}>0$), Fig. 3(e–f) still shows substantial $hc/e$ components, resembling the topological case in the upper panels. So that the capactiance based parity signal is not a definitive indicator for a topological phase.
@@
Figure 3: (a) Wave-function of the lowest energy eigenstate, (b) energy in K and (c) capacitance for the occupied and unoccupied state as a function of magnetic flux $\phi$ (in units where $hc/e=2$) for an $L=3\,\mu m$ Majorana nanowire in the topological patch regime [11] at $\mu=0.2\,K$ and $B_{x}=0.42\,T$, $V_{QD}=0.2\,K$, disorder strength $V_{RMS}=7.5\,K$. The wave-functions for this case show a single low energy ABS with a topological stability [1] (see Appendix A for details) $T_{S}=-0.621$. As expected [8] we see an approximate shift of $\Delta\phi=1$ between the even and odd parity sectors. The panels (d,e,f) are the corresponding result for the Majorana nanowire in a non-topological phase with $T_{S}=0.13$. The parameters for the phase are $B_{x}=0.56\,T$, $\mu=2.0\,K$, $V_{QD}=2.0\,K$ $V_{RMS}=15\,K$ and $L=3\,\mu m$.
II.2 Constraints on topology from parity oscillations
Comparing the results for our simulated flux and Fermion parity dependence of the capacitance for the ideal MZM case (i.e. Fig. 2(a,b,c)) to the other cases suggest that the capacitance in the topological case is characterized by an $hc/e-$periodic flux dependence together with an exact $hc/2e$ phase shift between the different parity sectors. Indeed, this signature is equivalent to both the fractional Josephson effect [9, 1, 2, 3] as well as teleportation [26, 17, 10, 3, 27] and thus can place complementary constraints on the topological character of a system compared to transport.
The fractional Josephson effect [9, 1, 2, 3] in the topological superconducting phase can be used to argue for a characteristic flux dependence of the capacitance. To understand this, note that for a ”long” superconducting wire, which is longer than the coherence length, the introduction of a phase slip in the middle of the superconducting wire is equivalent to an $hc/2e$ phase shift. In the trivial phase with a large length, this cannot have an effect on the capacitance, which must then be $hc/2e-$ periodic. In contrast, in the topologically non-trivial phase, such a phase slip must be accompanied by a change in fermion parity [1]. Thus, the capacitance in the topological case must be characterized by an $hc/2e-$ phase shift between the two fermion parity sectors, which must otherwise be $hc/e-$periodic. This argument assumes that the superconductor is infinitely long either in the topological or trivial phase. These conclusions are modified for a practical device where finite-size corrections, often strong need to be accounted for. A gapped trivial phase shows an $hc/2e-$periodic flux dependence similar to Fig. 2(d,e,f). The topological patch regime Fig. 3, while showing a near $hc/2e$ phase shift between the different parity sectors, also shows a change in average as well as amplitude that are not allowed in the topological phase. The recent capacitance measurements in the Majorana nanowire [8], do not distinguish between MZM and the topological patch regime because of an ambiguous flux periodicity.
The above argument, while suggesting the necessity of a specific flux-periodicity for the topological superconducting phase, does not show the flux periodicity to be a sufficient condition for topology. To understand this, let us consider a finite length superconducting nanowire with a single low-energy ABS that has weight at both ends of the nanowire. A topological superconducting nanowire with a pair of MZMs at the ends would be an example of such a state, as would the topological patch regime described above. To be more specific we can represent this low energy ABS or pair of MZMs in a wire by a fermion operator $c=(\hat{\gamma}_{R}+i\hat{\gamma}_{L})2^{-1/2}$. Let us represent the low energy fermion operator on the site $N=0$ by $d$. Introducing the flux dependent phase $\phi$ (in units so that $hc/2e=1$) on the right tunnel junction:
$H=\delta c^{\dagger}c+2^{1/2}t_{R}e^{i\pi\phi}(\gamma_{R}+\zeta_{R}\gamma_{L})d$
$+2^{-1/2}t_{L}d(\gamma_{L}+\zeta_{L}\gamma_{R})+h.c.,$ (2)
$H=\delta c^{\dagger}c+d[t_{R}e^{i\pi\phi}\{(1+i\zeta_{R})c^{\dagger}+(1-i\zeta_{R})c\}$
$+t_{L}\{(i+\zeta_{L})c^{\dagger}+(-i+\zeta_{L})c\}]+h.c.$ (3)
Note that choosing $\zeta_{R,L}$ to be complex allows $H$ to contain arbitrary complex normal and Andreev tunneling amplitude at each junction. The above Hamiltonian results from a projection of the BdG Hamiltonian in Eq. 9 to the lowest energy states in the nanowire and the QD. The operator $c$ corresponds to the lowest pair of states in the Majorana nanowire and the operator $d$ corresponds to the electron in the quantum dot. The parameter $\delta$ is directly controlled by the splitting energy of the MZMs or the voltage $V_{QD}$ of the quantum dot. The parameters $t_{L,R}$ and $\zeta_{L,R}$ are matrix elements of the tunneling term (proportional to $\lambda_{QD}$) in Eq. 9. Since the conclusions of this subsection depend on whether the relative order of magnitudes of these parameters, we do not compute them for the specific results in Fig. 2 and Fig. 4.
The two terms separate naturally by fermion parity sector. In the even fermion parity sector $c^{\dagger}c+d^{\dagger}d=0,2$, the $cd^{\dagger}=c^{\dagger}d=0$ so that writing $d^{\dagger}d=c^{\dagger}c=(1+\nu_{z})/2$ and $c^{\dagger}d^{\dagger}=\nu_{+}$ we get
$H_{even}(\phi)=\delta\nu_{z}/2$
+$+\{t_{R}e^{i\pi\phi}(1-i\zeta_{R})+t_{L}(-i+\zeta_{L})\}\nu_{-}+h.c.$ (4)
Here $\nu_{x,y,z}$ are $2\times 2$ Pauli matrices representing the fermion parity of the nanowire Fermion state $c^{\dagger}$. The matrices $\nu_{\pm}=(\nu_{x}\pm i\nu_{y})/2$. In the odd fermion parity sector $c^{\dagger}c+d^{\dagger}d=1$, the $c^{\dagger}d^{\dagger}=cd=0$ so that writing $21-d^{\dagger}d=c^{\dagger}c=(1+\nu_{z})/$ 2 and $dc^{\dagger}=\nu_{+}$ we get
$H_{odd}(\phi)=\delta\nu_{z}/2$
+$+\{t_{R}e^{i\pi\phi}(1+i\zeta_{R})+t_{L}(i+\zeta_{L})\}\nu_{+}+h.c.$ (5)
Writing the Hamiltonian in each case as $H=(\delta/2)\nu_{z}+A\nu_{+}+A^{*}\nu_{-}$ we can write the capacitance as
where the amplitude for the even and odd case is
$|A|^{2}$ $=|t_{R}e^{i\pi\phi}(1-i\zeta_{R})+t_{L}(-i+\zeta_{L})|^{2}\textrm{ for even parity}$
$=|t_{R}e^{i\pi\phi}(1+i\zeta_{R})+t_{L}(i+\zeta_{L})|^{2}\textrm{for odd parity}.$ (7)
The $\phi-$dependence of $|A|^{2}$ in both cases can be written in terms of three parameters $P,Q$ and $R$ as $|A|^{2}=P+Q\cos{\pi\phi}+R\sin{\pi\phi}$ ). In principle, Eqn. 6 together with this parametrization allows a fit for $\delta,P,Q$ and $R$ from the flux dependence of the experimentally measured capacitance $C(\phi)$ over one period. The parameters $P,Q,R$ for each parity sector combine to form six parameters, which can be combined using Eq. 7 to determine $t_{L,R}$ and the real and imaginary parts of $\zeta_{L,R}$. The central obstruction to performing this in the current experimental set-up [8] is the variation of the oscillation between different $hc/e$ periods, which indicates a magnetic-field induced background that must be removed.
Note that the MZM case with $\zeta_{L,R}=0$ is characterized by $P$ and $Q$ remaining unchanged between the two parity sectors while $R$ changes sign. This corresponds to a $hc/2e$ shift of $\phi$, which is the essential signature expected for a topological phase. On the other hand, MZM also requires that $\delta$ is set entirely by the QD voltage off-set $V_{QD}$, whose estimation is subject to thermal noise. Thus, the extent to which the capacitance oscillations show an $hc/2e$ parity shift can bound the spurious coupling of the MZMs though it is not a sensitive test of the vanishing of the MZM splitting energy.
@@ III Rabi oscillations
where $H_{L,R}$ are given by Eq. 9 except that the sites are restricted to $r\neq 0$ together with $\lambda_{QD}=0$ to remove the QD and $\lambda(\tau)$ is the dimensionless tunneling strength between the wires. Note that this set-up differs fundamentally from previous proposals for studying coherence in nanowire qubits [28], which requires inter-wire parity measurements. Instead, the set-up in Fig. 1, similar to previous experiments [13], is based on intra-wire parity measurements together with inter-wire tunneling that is set to a non-zero value $\lambda(\tau>0)=\lambda$ after initializing with $\lambda(\tau=0)=0$.
In this set-up for Rabi oscillations one would initialize the system at $\lambda(\tau=0)=0$ where each nanowire has a single low energy quasiparticle whose occupation is determined by the initial parity and then increase the tunneling to $\lambda(\tau>0)=\lambda$ to generate Rabi oscillations for the quasiparticle to be in the left |$\ket{L}$ or right wire |$\ket{R}$. This initializes the system in a superposition of the four lowest energy eigenstates of the final BdG Hamiltonian. The four eigenstates can be classified into groups of two based on the total fermion parity sector. Restricting the dynamics to the lowest two positive energy state (one from each nanowire) and considering a specific total Fermion parity sector allows one to map the dynamics using the Bloch-Redfield master equation of a standard two-level spin system [29] that is characterized by a Rabi oscillation frequency $\Omega$, relaxation time $T_{1}$, dephasing time $T_{2}$ as well as an amplitude for the Rabi oscillation of the Fermion parity of the left wire (see Appendix B for details).
Note that the set-up of the Majorana qubit shown in Fig. 1 and the qubit degree of freedom described above is essentially similar to a charge qubit [30, 31, 32]. The dominant source of decoherence in this case is charge noise from Coulomb impurities in the environment, which tends to be long-range correlated. Thus, we will approximate this charge noise as an electrostatic potential that is Markovian i.e. uncorrelated in time but uniform across the length of each wire. The baths of each wire are assumed to be uncorrelated with each other. In the BdG formalism, such a noise Hamiltonian corresponds to uncorrelated in time fluctuations of the chemical potential parameter $\mu$ in Eq. 9(in Appendix A). The dynamics of the density matrix generated by the Hamiltonian $H_{tot}$ as well as coupling to baths is described by a Bloch-Redfield master equation [29] (see Appendix C for details). The effect of the charge noise bath in the wires $L$ and $R$ are represented by parameters $\gamma_{L,R}$ in the master equation 11 in Appendix C. The magnitude of the bath coupling parameters are set to $\gamma_{L,R}\sim 80\,mK$ so as to reproduce conventional charge qubit dephasing [31].
In addition to Rabi oscillations, we will also consider the possibility of estimating the coherence time of Majorana qubits [28] using inter-wire parity or capacitance measurements instead of the inter-wire tunneling required for Rabi oscillations. Such inter-wire parity measurements are necessary to perform topological operations [6]. The estimate of dephasing rates in this case takes the form of a decoherence time $T_{1}^{\prime}$ for the inter-wire fermion parity. Different choices of ends for the inter-wire fermion parity measurement $X$ and $Y$ are expected to generate different superpositions that differ by a phase shift $\delta\phi$ that can also characterize a topological qubit (see Appendix D for details).
III.1 Coherence results